OpenStudy (anonymous):

Polygon ABCD has the following vertices: A(−5, 4), B(1, 4), C(6, −4), and D(−5, −4) Calculate the area of the polygon. 48 units squared 53 units squared 68 units squared 88 units squared

OpenStudy (anonymous):

help pleez

OpenStudy (anonymous):

its B

OpenStudy (anonymous):

k ill try but i could be wrong

OpenStudy (anonymous):

will give medal and be fan

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

explane pleez?

OpenStudy (anonymous):

@I_Always_Smiling

OpenStudy (anonymous):

@sammixboo

OpenStudy (sleepyjess):

@iamwavy , Welcome to OpenStudy! Please do not give out direct answers. @help100questions , are you familiar with the distance formula?

OpenStudy (anonymous):

no

OpenStudy (sleepyjess):

Ok, well it is essential here. The distance formula is \(\sqrt{(x_2-x_1)^2 + (y_2-y_1)^2}\). So just pick 2 points and I will show you how to use this.

OpenStudy (anonymous):

-5, 4 and 1,4

OpenStudy (sleepyjess):

Ok, so -5, 4 is our \(x_1,y_1\) and 1, 4 is \(x_2,y_2\).

OpenStudy (anonymous):

ok

OpenStudy (sleepyjess):

Now just substitute the values into the formula. \(\sqrt{(1+5)^2 + (4-4)^2}\)

OpenStudy (anonymous):

so -5 - 4 = 9 and 1 - 4 = 3 so thats confusing

OpenStudy (sleepyjess):

I gave you all the substitutions.

OpenStudy (anonymous):

the nubera in the parenthese are subtratred?

OpenStudy (anonymous):

sorry for my bad spelling

OpenStudy (sleepyjess):

yes, but subtracting a negative number is the same as adding

OpenStudy (anonymous):

oh

OpenStudy (sleepyjess):

it's fine, please excuse any typos I have, I was up until 5 am

OpenStudy (anonymous):

so 9 + 3 = 12

OpenStudy (sleepyjess):

This is after substituting the numbers into the formula \(\sqrt{(1+5)^2 + (4-4)^2}\)

OpenStudy (anonymous):

wait 6 + 0?

OpenStudy (sleepyjess):

No, \(\sqrt{6^2 + 0^2}\)

OpenStudy (anonymous):

and thats 6 soooooo

OpenStudy (sleepyjess):

what is \(6^2\)?

OpenStudy (anonymous):

ummm 60?

OpenStudy (sleepyjess):

6*6

OpenStudy (anonymous):

or 12? im bad at math

OpenStudy (anonymous):

36 ok

OpenStudy (sleepyjess):

\(6^2\) is the same as \(6*6\)

OpenStudy (anonymous):

36

OpenStudy (anonymous):

SRY!!!!!!

OpenStudy (sleepyjess):

it's ok :)

OpenStudy (anonymous):

<3

OpenStudy (sleepyjess):

ok so now we have \(\sqrt{36}\) which we already know is 6

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

thanks

OpenStudy (sleepyjess):

That isn't all

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

???????

OpenStudy (anonymous):